# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v4

@@ 2.2.1 PnP prior
Venkatakrishnan et al. proposed the first PnP method based on ADMM [56]. PnP methods replace the proximal mapping of the implicit regularizer defined by (9) with a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. PnP-ADMM (Algorithm 1) is well known for its fast empirical convergence and efficiency in computational imaging [35, 60]. Sreehari et al. established theoretical conditions for PnP-ADMM, requiring that the Jacobian $\nabla D_{\sigma}$ be a doubly stochastic and symmetric matrix with all real eigenvalues in the range $(0,1]$ [51]. When the denoiser meets the necessary and sufficient conditions of Proposition 2.5, it is a proximal mapping of the implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$. Subsequently, a non-local means denoiser satisfying these conditions was constructed [51].
@@
Given $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the proximal operator of $\phi(\mathbf{x})$ is defined by
(9) $\mathrm{prox}_{\phi}(\mathbf{x})=\arg\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}+\phi(\mathbf{x}).$
@@
A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex l.s.c. (lower semicontinuous) function $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ if, and only if the following conditions hold jointly:
@@
    there exists a convex l.s.c. function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
@@
    $\left\|h(\mathbf{x})-h(\mathbf{y})\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}.$
@@ 2.2.2 RED prior
Romano et al. introduced a famous RED [47], which is defined by
@@
where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser which is assumed to obey the following assumptions,
@@
    Jacobian Symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x}),\forall\mathbf{x}\in\mathbb{R}^{n}$.
@@ 2.2.3 Implicit Gradient RED
The gradient of RED is exactly the denoising residual. We show that under ideal conditions and assuming the residual is nonexpansive, both the denoiser and its residual function as proximal operators of implicit regularizers. Even if all RED conditions are hard to satisfy, learning FNE operators can still achieve this.
The nonexpansive assumption has theoretically played a crucial role in PnP and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49],
(A) $\left\|(I-D_{\sigma})(\mathbf{u})-(I-D_{\sigma})(\mathbf{v})\right\|\leq\varepsilon\left\|\mathbf{u}-\mathbf{v}\right\|,$
for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ and $\varepsilon\leq 1$, the real spectral normalization is used to obtain nonexpansive residual. Under the conditons C1C2C3 and C4 of RED and assumption (A), we can further discuss the relationship between $\nabla g_{\text{red}}$ and the proximal operator, and prove that there exists an implicit regularizer $\phi$ such that $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$ and $R(\mathbf{x})=\nabla g_{\text{red}}(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$.
@@
Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies conditons C1C2C3C4 and assumption (A), then there exists an implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ such that
@@
By assumption condtions, the RED term $g_{\text{red}}$ is convex and $\nabla g_{\text{red}}=I-D_{\sigma}=R$. Let $\psi(\mathbf{x})=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$, then $D_{\sigma}=I-\nabla g_{\text{red}}=\nabla\psi$, which is a GS denoiser. Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex. Applying Proposition 2.5 together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=I$, then there exist a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
@@
Moreover, by [2, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$. Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator. Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudo-contractive denoisers [57], have made the nonexpansive assumption more realistic in practice. Constructing the FNE denoiser $D_{\sigma}$ is feasible, and similarly, the residual $R=I-D_{\sigma}$ can also be FNE in (7) with $\mathcal{M}=I$. They can both be the proximal operators of implicit regularizers.
@@ 3 Halpern-type preconditioned proximal point (HPPP)
Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29]. The Halpern iteration [26]
$\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$
is an implicitly regularized method to find a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$. When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$. Based on the Halpern iteration, we propose the Halpern-type preconditioned proximal point (HPPP) algorithms (5).
@@ 3.1 Convergence analysis
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5). Assume that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies
@@
By Lemma A.2(i), the sequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ is bounded, there exists a weakly convergent subsequence $\{\mathbf{u}^{k_{n_{j}}}\}_{j\in\mathbb{N}}$, we may assume $\mathbf{u}^{k_{n_{j}}}\rightharpoonup\tilde{\mathbf{u}}$. According to the conditions of the theorem, we have $\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$.
Applying Lemma A.2 together with conditions (i)-(iii), we obtain $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to\mathbf{0}$, then $\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\rangle_{\mathcal{M}}\to 0$. Hence, we have
@@
Denote $\delta_{k+1}=\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$, notice that
@@
applying Lemma A.1 together with conditions (i)-(ii) and $\limsup_{k\to\infty}\delta_{k}\leq 0$, then we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0.$
@@
where the last inequality follows from (26) and $C$ is defined in (25). Since $\lim_{k\to\infty}\mu_{k}=0$ and $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$, and thus combined with (14) yields that
@@
It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfes conditions (i)-(iii) of Theorem 3.1. Compared with [3], we have provided an alternative method to obtain the limit point as $\mathcal{M}$-projection of the initial point. The work [3] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$. By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. According to Theorem 3.1, we obtain the following Proposition 3.3.
@@
Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. By Lemma A.2(iii), we have $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, it follows that
@@
i.e., $\mathcal{T}\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. Using $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$ we have
@@
Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$. Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. This completes the proof of Proposition 3.3.
@@
Compared Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$. All conditions are the same except for additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$. The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption especially in applications to splitting algorithms, which is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly, where $U$ is a closed convex cone and $V$ is a closed hyperplane [31]. Building on Hundal’s counterexample, [13] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence. Similarly, using the same counterexample, we can demonstrate that PPP also fails to converge strongly.
@@
$\mathcal{A}_{1}:\mathbf{x}\mapsto\begin{cases}V^{\perp}&\text{ if }\mathbf{x}\in V,\\ \emptyset&\text{ if }\mathbf{x}\in V,\end{cases}\text{ and }\mathcal{A}_{2}=(P_{V}\circ P_{U}\circ P_{V})^{-1}-I.$
@@
As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the following sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
(15) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+J_{\mathcal{A}_{2}}(2J_{\mathcal{A}_{1}}\mathbf{w}^{k}-\mathbf{w}^{k})-J_{\mathcal{A}_{1}}\mathbf{w}^{k}$
generated by the DRS algortihm can weakly converges to $\mathbf{0}$ and $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$.
Since the above (15) can be viewed as a special PPP with
$\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&\mathcal{A}_{2}\end{pmatrix},\mathcal{M}=\begin{pmatrix}I&-I\\ -I&I\end{pmatrix},$
and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$. By [9, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$. Indeed, if by contraction $\mathbf{u}^{k}\to\mathbf{0}$, then
@@
which contradicts the fact $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$. Thus $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can only converge weakly to $\mathbf{0}$.

### v5

@@ 2.2.1 PnP prior
Venkatakrishnan, Bouman, and Wohlberg proposed the first PnP method based on ADMM [56]. PnP methods replace the proximal mapping of the implicit regularizer defined by (9) with a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. PnP-ADMM (Algorithm 1) is well known for its fast empirical convergence and efficiency in computational imaging [35, 60]. Sreehari et al. established theoretical conditions for PnP-ADMM, requiring that the Jacobian $\nabla D_{\sigma}$ be a doubly stochastic and symmetric matrix with all real eigenvalues in the range $(0,1]$ [51]. When the denoiser meets the necessary and sufficient conditions of Proposition 2.5, it is a proximal mapping of the implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$. Subsequently, a nonlocal means denoiser satisfying these conditions was constructed [51].
@@
The proximal operator of $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is defined by
(9) $\mathrm{prox}_{\phi}(\mathbf{x})=\arg\min_{\mathbf{u}\in\mathbb{R}^{n}}\{\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}+\phi(\mathbf{u})\}.$
@@
A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex lower semicontinuous function $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ if and only if the following conditions hold jointly:
@@
    There exists a convex lower semicontinuous function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
@@
    $\left\|h(\mathbf{x})-h(\mathbf{y})\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|\quad\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}.$
@@ 2.2.2 RED prior
RED [47] is defined by
@@
where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser which is assumed to obey the following assumptions:
@@
    Jacobian symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x})$ for all $\mathbf{x}\in\mathbb{R}^{n}$.
@@ 2.2.3 Implicit Gradient RED
The gradient of RED is exactly the denoising residual. We show that under ideal conditions and assuming the residual is nonexpansive, both the denoiser and its residual function are proximal operators of implicit regularizers. Even if all RED conditions are hard to satisfy, learning FNE operators can still achieve this.
The nonexpansive assumption has theoretically played a crucial role in PnP prior and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49]:
(A) $\left\|(I-D_{\sigma})(\mathbf{u})-(I-D_{\sigma})(\mathbf{v})\right\|\leq\varepsilon\left\|\mathbf{u}-\mathbf{v}\right\|;$
for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ and $\varepsilon\leq 1$, the real spectral normalization is used to obtain nonexpansive residual. Under the conditions C1C2C3, and C4 of RED and assumption (A), we can further discuss the relationship between $\nabla g_{\text{red}}$ and the proximal operator, and we prove that there exists an implicit regularizer $\phi$ such that $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$ and $R(\mathbf{x})=\nabla g_{\text{red}}(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$.
@@
Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies conditions C1C2C3C4 and assumption (A); then there exists an implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ such that
@@
By assumption conditions, the RED term $g_{\text{red}}$ is convex, and $\nabla g_{\text{red}}=I-D_{\sigma}=R$. Let $\psi(\mathbf{x})=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$, then $D_{\sigma}=I-\nabla g_{\text{red}}=\nabla\psi$, which is a GS denoiser. Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex. Applying Proposition 2.5, together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=\mathbf{x}$, there exists a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
@@
Moreover, by [2, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$. Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator. Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudocontractive denoisers [57], have made the nonexpansive assumption more realistic in practice. Constructing the FNE denoiser $D_{\sigma}$ is feasible, similarly, the residual $R=I-D_{\sigma}$ can be FNE in (7) with $\mathcal{M}=I$. Both can be proximal operators of implicit regularizers.
@@ 3 Halpern-type preconditioned proximal point (HPPP)
Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29]. The Halpern iteration $\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$ [26] is an implicitly regularized method for finding a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$. When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$. Based on the Halpern iteration, we propose the Halpern-type preconditioned proximal point (HPPP) algorithms (5).
@@ 3.1 Convergence analysis
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipschitz. Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5). Assume that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, and that $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies
@@
By Lemma A.2(i), the sequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ is bounded, there exists a weakly convergent subsequence $\{\mathbf{u}^{k_{n_{j}}}\}_{j\in\mathbb{N}}$, and we may assume $\mathbf{u}^{k_{n_{j}}}\rightharpoonup\tilde{\mathbf{u}}$. According to the conditions of the theorem, we have $\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$.
Applying Lemma A.2, together with conditions (i)-(iii), we obtain $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to\mathbf{0}$; then $\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\rangle_{\mathcal{M}}\to 0$. Hence, we have
@@
Denoting $\delta_{k+1}=\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$, noticing that
@@
and applying Lemma A.1, together with conditions (i)-(ii) and $\limsup_{k\to\infty}\delta_{k}\leq 0$, we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0.$
@@
where the last inequality follows from (26) and $C$ is defined as in (25). Since $\lim_{k\to\infty}\mu_{k}=0$ and $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$, combined with (14) they yield that
@@
It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfies conditions (i)-(iii) of Theorem 3.1. Compared with [3], we have provided an alternative method to obtain the limit point as the $\mathcal{M}$-projection of the initial point. The work [3] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$. By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. According to Theorem 3.1, we obtain the following proposition.
@@
Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. By Lemma A.2(iii), we have $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, and it follows that
@@
i.e., $\mathcal{T}\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. Using $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, we have
@@
Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], and hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$. Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. This completes the proof of Proposition 3.3.
@@
We compare Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$. All conditions are the same, except for the additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$. The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption, especially in applications to splitting algorithms, and is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly [31], where $U$ is a closed convex cone and $V$ is a closed hyperplane. Building on Hundal’s counterexample, [13] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence. Similarly, using the same counterexample, we can demonstrate that PPP also fails to converge strongly.
@@
$\mathcal{A}_{1}:\mathbf{x}\mapsto\begin{cases}V^{\perp}&\text{ if }\mathbf{x}\in V,\\ \emptyset&\text{ if }\mathbf{x}\notin V,\end{cases}\text{ and }\mathcal{A}_{2}=(P_{V}\circ P_{U}\circ P_{V})^{-1}-I.$
@@
As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
(15) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+J_{\mathcal{A}_{2}}(2J_{\mathcal{A}_{1}}\mathbf{w}^{k}-\mathbf{w}^{k})-J_{\mathcal{A}_{1}}\mathbf{w}^{k},$
generated by the DRS algorithm can weakly converge to $\mathbf{0}$ and $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$.
The sequence (15) can be viewed as a special PPP with
$\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&\mathcal{A}_{2}\end{pmatrix},\quad\mathcal{M}=\begin{pmatrix}I&-I\\ -I&I\end{pmatrix},$
and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$. By [9, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$. Indeed, if by contraction, $\mathbf{u}^{k}\to\mathbf{0}$, then
@@
which contradicts the fact that $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$. Thus $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can only converge weakly to $\mathbf{0}$.